# Impact of the carbon price on credit portfolio’s loss with stochastic collateral

### v1

@@ 2.1 A Multisectoral Model with Carbon price
In our framework, a representative firm in each sector which maximizes its profits by choosing, at each time and for a given productivity, the quantities of labor and intermediary inputs, while, a representative household solves a dynamic optimization problem to decide how to allocate its consumption expenditures among the different goods and hours worked and among the different sectors. We assume that the utility function $U(x,y):=\log{x}-\frac{y^{1+\varphi}}{1+\varphi}$ with $\varphi\geq 0$. Moreover, $\boldsymbol{\lambda}$ (respectively $\psi$) are matrix in $(\mathbb{R}_{+}^{*})^{I\times I}$ (respectively vector in $(\mathbb{R}_{+}^{*})^{I}$) of the elasticities of intermediary inputs (respectively labor). We also assume a constant return to scale, namely
@@
Since the productivity and the carbon price processes are continuous, the firms and households problems are well posed and their solutions exist. More details are given in B. In the following proposition, we give the expression of the output.
@@
with $\overline{\mathfrak{d}}:=(\overline{\tau}\overline{\delta},\overline{\zeta}\overline{\delta},\overline{\kappa}\overline{\delta})$. Assume that
@@
    $\mathfrak{e}_{t}=\mathfrak{e}(\mathfrak{d}_{t}):=(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top})^{-1}\mathbf{1},$ (2.15)
@@
    Using $\mathcal{B}_{t}=(\mathcal{B}^{i}_{t})_{i\in\mathcal{I}}:=\left[\mathcal{A}_{t}^{i}+v^{i}(\mathfrak{d}_{t})\right]_{i\in\mathcal{I}}$ with for $(\overline{\tau},\overline{\zeta},\overline{\kappa},\overline{\delta})\in\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}$,
@@
    ${Y}_{t}=\exp{\left((\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}{\mathcal{B}_{t}}\right)}.$ (2.17)
@@
    Furthermore, since $\mathfrak{d}\in\mathcal{C}^{1}(\mathbb{R}_{+},[0,1)^{I}\times(\mathbb{R}_{+})^{I\times I}\times(\mathbb{R}_{+})^{I})$, we directly have $\Psi(\overline{\mathfrak{d}}_{\cdot}),\Lambda(\overline{\mathfrak{d}}_{\cdot})\in\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R})$. Moreover, $\bar{\mathfrak{d}}\mapsto(\mathbf{I}_{I}-{\Lambda}(\bar{\mathfrak{d}})^{\top})^{-1}$ on $\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}$) is differentiable, then $(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{\cdot})^{\top})^{-1})\in\mathcal{C}^{1}(\mathbb{R}_{+},\mathbb{R})$.
The output ${Y}$ is also positive, we can then introduce, from the third item, for all $t\geq 0$, $\mathrm{d}\log{({Y}_{t})}$ representing the instantaneous consumption growth. This proposition is an equivalent of Bouveret et al., (2023)[Theorem 1.10.] in continuous time.
@@ 2.2 A Firm Valuation Model
Consider a fix $n\in\{1,\ldots,N\}$. For any time $t\in\mathbb{R}_{+}$ and firm $n$, we note $F^{n}_{t}$ the free cash flows of $n$ at $t$, and $r>0$ the discount rate, we introduce the following assumption:
@@
We define the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ by $\mathcal{F}_{t}=\sigma\left(\mathcal{G}_{t}\cup\sigma\left\{\mathfrak{b}_{s}:s\in[0,t]\right\}\right)$ for $t\geq 0$, and we denote $\mathbb{E}_{t}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{t}]$.
@@
For $n\in\{1,\ldots,N\}$ and $t\geq 0$, describing $V^{n}_{t,\mathfrak{d}}$ as a function of the underlying processes driving the economy does not lead to an easily tractable formula. To facilitate the forthcoming credit risk analysis, when $\varsigma$ (introduced in Standing Assumption 2.1) is closed to 00, we approach $V^{n}_{t,\mathfrak{d}}$ by the quantity
@@
that we describe as a proxy the firm $n$ value at time $t$. We will work directly with $\mathcal{V}^{n}_{t,\mathfrak{d}}$ instead of $V^{n}_{t,\mathfrak{d}}$. We have the following proposition, whose proof is given in B.4.
@@
    Moreover, with $t_{\circ}$ and $t_{\star}$ defined in Standing Assumption 2.3, we obtain the following explicit form,
@@
    $\rho_{n}:=\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}\lVert\mathfrak{a}^{n\cdot}\rVert^{2}\lVert\Sigma\rVert^{2}<r.$ (2.24)
@@
The following corollary gives (conditional) laws of the (proxy) of the firm value $\mathcal{V}_{\cdot,\mathfrak{d}}^{n}$.
@@ 2.3 A Credit Risk Model without collateral
To conclude this section, we present the probability of default. As Basel Committee on Banking Supervision, (2017), we introduce four credit risk parameters: the probability of default (PD) measures the default risk associated with each borrower, the exposure at default (EAD) measures the outstanding debt at the time of default, the loss given default (LGD) denotes the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity $T$ represents the duration of the credit. With these four parameters, we can compute the portfolio loss $L$, with some assumptions:
@@
We follow in the present work the second definition of default proposed: a firm default when it is over-indebted, that is in fact the same approach used in the structural credit risk models. We retain this definition in this continuous setting in the same way as Bourgey et al., (2022). Actually, the term default may even be considered an abuse of language. So, to avoid any confusion in the following, we will use the terms default and over-indebted without distinction. Therefore, the over-indebtedness of entity $n$ occurs at time $t$ when the firm value $\mathcal{V}^{n}_{t,\mathfrak{d}}$ falls below a given barrier $\mathcal{D}_{t}^{n}$, related to the net debt, namely on the event $\left\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}$.
However, it should be noted that in a continuous time setting, it can be interesting to work in the Black and Cox, (1976) model. Here, the default event depends on the trajectory of the firm value process $\mathcal{V}$. Therefore, at a given time $t$, the firm defaults if it has been over-indebted at least one time during the period $[0,t]$, that is $\left\{\exists s\in[0,t]\text{ such that }\mathcal{V}^{n}_{s,\mathfrak{d}}<\mathcal{D}^{n}_{s}\right\}$. Thus, the default time is given by
$\tau^{n}:=\inf\left\{t\geq 0,\quad\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}.$ (2.27)
@@
therefore $\log{\mathcal{V}^{n}_{\cdot,\mathfrak{d}}}$ is a Gaussian process. However, as Azaıs and Wschebor, (2000) summarizes, the computation of the distribution function of the random variable $\inf_{0\leq s\leq t}\log{\mathcal{V}^{n}_{s,\mathfrak{d}}}$ is by means of a closed formula is known only for a very restricted number of stochastic processes as the Brownian Motion, the Brownian Bridge, the Brownian Motion with a linear drift, and the stationary Gaussian processes with relatively simple with covariance. This is not the case here.
@@
where $\mathcal{K}^{n}(\mathfrak{d},t,T,a,\theta)$ and $\mathcal{L}^{n}(t,T)$ are defined in Corollary 2.7.
The previous results tell us that the probability of the over-indebtedness depends on:
@@
    parameters specific to the climate transition
@@
    parameters specific to the company (the contract),
@@
    parameters specific to the economy to which the company belongs to:
@@
Therefore, by assuming that EAD and LGD are deterministic and independent of the carbon price, we could obtain the expressions of the EL and UL. In the next section, we will express LGD as a function of some guarantees which are affected by the climate transition.
@@ 3 LGD with stochastic collaterals in continuous time
We are in the same framework as in the previous section, but Assumption 2.8(2) is not satisfied anymore, therefore the bank could require from each counterpart $1\leq n\leq N$ a (single) collateral $C^{n}$ to secure its debt. Collateral can take the form of a physical asset such as real estate, business equipment, or inventory, or it can be a financial asset such as invoices, cash, or investments. If a firm is over-indebted at time $t$, we assume that the liquidation ends at $t+a$ with $a\in\mathbb{R}_{+}$ where $a$ is the liquidation delay. Moreover, $k\in[0,1)$ represents the fraction of the collateral used to cover liquidations auctions, as well as other legal and administrative procedures.
@@
The potential loss that would be recorded due to the firm default event is the difference between the debt amount $\mathrm{EAD}_{t}^{n}$ and the amount gets after the recovery processes at the time horizon. Consequently, if there is not default ($\mathcal{V}^{n}_{t,\mathfrak{d}}\geq\mathcal{D}^{n}_{t}$) or there is default and the collateral liquidated exceed the exposure ($(1-k)e^{-ra}C^{n}_{t+a}\geq\mathrm{EAD}_{t}^{n}$), the loss noted $L_{n,t}$ is zero, and if there is default and if the exposure exceed the collateral liquidated, the loss is
$L_{n,t}=\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}-\gamma(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})=(1-\gamma)\left(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}\right),$ (3.1)
where the constant $r$ is the discount rate. The loss of the portfolio at time $t$, is in fact, defined as
$L^{N}_{t}:=\sum_{n=1}^{N}L_{n,t}=\sum_{n=1}^{N}(1-\gamma)(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}.$ (3.2)
@@
$\mathrm{L}^{\mathbb{G},N}_{t}:=\sum_{n=1}^{N}L_{n,t}^{\mathbb{G}}\quad\text{with}\quad L_{n,t}^{\mathbb{G}}=\mathbb{E}\left[L_{n,t}\middle|\mathcal{G}_{t}\right]=\mathrm{EAD}_{t}^{n}\cdot\mathrm{LGD}_{t,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t,\mathfrak{d}}^{n},$ (3.3)
@@
Explicitly, in the above theorem, we assume that our portfolio is perfectly fine grained, so that we can approximate $L_{t}^{N}$ – the portfolio loss – by $L_{t}^{\mathbb{G},N}$ – the conditional expectation of loss given the systemic factor–. By construction, the loss given default noted $\mathrm{LGD}$ is the percentage of the total exposure that is lost when a over-indebtedness occurs. The literature on LGD modeling is fairly extensive. We can distinguish namely, economic modeling Bastos, (2010); Roncalli, (2020) and stochastic modeling Roncalli, (2020)[Page 193], Chalupka and Kopecsni, (2008). As the definition of $\mathrm{PD}$ does not change compare to what we did in Section 2 and as $\mathrm{EAD}$ is given, we will focus on $\mathrm{LGD}$ modeling. We can first remark that $0\leq\mathrm{LGD}_{t,\mathfrak{d}}^{n}\leq 1-\gamma$, then the presence of a collateral necessarily reduces $\mathrm{LGD}$.
Key quantities for the bank to understand the (dynamics of the) risk in the portfolio are the (expected and unexpected) losses and probability of default conditionally to the (information generated by the) risk factors. Precisely, for a date $t$ and a horizon $T$, a bank would like to know some risk measures at $t$ of its portfolio maturing at horizon $T$.
@@
    The (conditional) Expected Loss (EL) reads
@@
    The Unexpected Loss (UL) reads for $\alpha\in(0,1)$,
    $\mathrm{UL}^{N,T}_{t}\!(\alpha):=\mathrm{VaR}^{\alpha,N,T}_{t}-\mathrm{EL}^{N,T}_{t},\quad\text{where}\quad 1-\alpha=\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right].$ (3.6)
@@
    The Stressed Loss (or Expected Shortfall or ES) reads $\mathrm{VaR}^{\alpha}_{t}(L^{N}_{s})$:
@@
From now, if the collateral exists, we focus on two types: a financial asset and a property in housing market. We introduce the two following lemmas which will help later on to set up explicit formulas for LGD. Let $n\in\{1,\ldots,N\}$.
@@
$\rho_{t,T,a}^{n}:=\frac{cv_{t,T,a}^{n}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}},$
and
$\overline{\omega}_{t,T,a}^{n}:=\frac{\log{\left(u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\right)}-\overline{\mathcal{K}}^{n}_{\mathfrak{d},t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}}.$
@@
However, according to (A.3), $\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]=e^{\frac{1}{2}\sigma^{2}}\Phi_{2}\left(x-\sigma,y-\rho\sigma;\rho\right)$, therefore,

### v2

@@ 3.1 A Multisectoral Model with Carbon price
In our framework, a representative firm in each sector which maximizes its profits by choosing, at each time and for a given productivity, the quantities of labor and intermediary inputs, while, a representative household solves a dynamic optimization problem to decide how to allocate its consumption expenditures among the different goods and hours worked and among the different sectors. We assume that the utility function $U:\mathbb{R}_{+}^{*}\times\mathbb{R}_{+}^{*}\to\mathbb{R},(c,h)\mapsto\log{c}-\frac{h^{1+\varphi}}{1+\varphi}$ with $\varphi\geq 0$; $c$ and $h$ represents respectively the household consumption and the labor. Moreover, $\boldsymbol{\lambda}$ (respectively, $\psi$) are matrix in $(\mathbb{R}_{+}^{*})^{I\times I}$ (respectively vector in $(\mathbb{R}_{+}^{*})^{I}$) of the elasticities of intermediary inputs (respectively labor). We also assume a constant return to scale, namely
@@
Since the productivity and the carbon price processes are continuous, the firms and households problems are well posed and their solutions exist. More details are given in appendix B. In the following proposition, we give an explicit expression of the output.
@@
with $\overline{\mathfrak{d}}:=(\overline{\delta}\overline{\tau},\overline{\delta}\overline{\zeta},\overline{\delta}\overline{\kappa})$. Assume that
@@
    $\mathfrak{e}_{t}=\mathfrak{e}(\mathfrak{d}_{t}):=(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{t})^{\top})^{-1}\mathbf{1}.$ (3.12)
@@
    For $(\overline{\tau},\overline{\zeta},\overline{\kappa},\overline{\delta})\in\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}$,
@@
    ${Y}_{t}=\exp{\left((\mathbf{I}_{I}-\boldsymbol{\lambda})^{-1}{\left(\mathcal{A}_{t}+v(\mathfrak{d}_{t})\right)}\right)}.$ (3.14)
@@
    Furthermore, since $\mathfrak{d}\in\mathfrak{D}(\mathbb{R}_{+},[0,1)^{I}\times(\mathbb{R}_{+})^{I\times I}\times(\mathbb{R}_{+})^{I})$, we directly have $\Psi(\overline{\mathfrak{d}}_{\cdot}),\Lambda(\overline{\mathfrak{d}}_{\cdot})\in\mathfrak{D}(\mathbb{R}_{+},\mathbb{R})$. Moreover, $\bar{\mathfrak{d}}\mapsto(\mathbf{I}_{I}-{\Lambda}(\bar{\mathfrak{d}})^{\top})^{-1}$ on $\mathbb{R}_{+}^{I}\times\mathbb{R}_{+}^{I\times I}\times\mathbb{R}_{+}^{I}$) is differentiable, then $(\mathbf{I}_{I}-{\Lambda}(\mathfrak{d}_{\cdot})^{\top})^{-1})\in\mathfrak{D}(\mathbb{R}_{+},\mathbb{R})$.
The output ${Y}$ is also positive, we can then introduce, from the third item, for all $t\geq 0$, $\mathrm{d}\log{({Y}_{t})}$ representing the instantaneous consumption growth. This proposition is an equivalent of Bouveret et al., (2025)[Theorem 1.10.] in continuous time. In the same spirit, we can also define and determine
  • 1.
    consumption, intermediary inputs, labor, and GHG emissions at each time as well as their logarithmic growth between two instants.
  • 2.
    The sensitivities of the growth of variables to the carbon price. This allows us to answer the following types of questions: what is the impact on all variables if in the whole economy, we only charge GHG emissions coming from the production of companies in sector $1$? The main difficulty here comes from the fact that the coefficients of the inverse of the elasticity matrix $(\boldsymbol{\lambda}^{ji})_{i,j\in\mathcal{I}}$ are not known.
  • 3.
    The law of the logarithmic growth of macroeconomic variables that we obtain by using the Gaussianity and stationarity of logarithmic growth.
  • 4.
    the evolution of the contribution of each sector in production/consumption as well as the evolution of GHG emissions.
@@ 3.2 A Firm Valuation Model
The goal in this section is to describe the dynamics of the value of a firm operating in an economy that undergoes the climate transition described by both the carbon price and carbon intensities.
Consider a firm $n\in\{1,\ldots,N\}$. For any time $t\in\mathbb{R}_{+}$ and firm $n$, we note $F^{n}_{t}$ the free cash flows of $n$ at $t$, and $r>0$ the discount rate, we introduce the following assumption:
@@
We call $\tilde{\mathfrak{a}}^{n\cdot}=(\tilde{\mathfrak{a}}^{ni})_{1\leq i\leq I}$ (respectively $\mathfrak{a}^{n\cdot}=(\mathfrak{a}^{ni})_{1\leq i\leq I}$) factor loadings, quantifying the extent to which $\mathrm{d}\log{F^{n}_{,\mathfrak{d}}}$ is related to $\mathrm{d}\log{{Y}}$ (respectively $\mathrm{d}\mathcal{A}$). We define the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ by $\mathcal{F}_{t}=\sigma\left(\mathcal{G}_{t}\cup\sigma\left\{\mathfrak{b}_{s}:s\in[0,t]\right\}\right)$ for $t\geq 0$, and we denote $\mathbb{E}_{t}[\cdot]:=\mathbb{E}[\cdot|\mathcal{F}_{t}]$.
@@
For $n\in\{1,\ldots,N\}$ and $t\geq 0$, describing $V^{n}_{t,\mathfrak{d}}$ as a function of the underlying processes driving the economy does not lead to an easily tractable formula. But this allows us to write it as a fixed-point problem which can be solved by numerical methods such as deep learning methods (see Hammad et al., (2022); Kobeissi and Bach, (2022)) or Picard iteration (see Berinde and Takens, (2007)). However, to facilitate the forthcoming credit risk analysis, when $\varsigma$ (introduced in 3.1) is close to 00, we approach $V^{n}_{t,\mathfrak{d}}$ by the quantity
@@
which is defined using the first order expansion of $\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}$ and that we describe as a proxy the firm $n$ value at time $t$. We will work directly with $\mathcal{V}^{n}_{t,\mathfrak{d}}$ instead of $V^{n}_{t,\mathfrak{d}}$. We have the following proposition, whose proof is given in B.4.
@@
    Moreover, with $t_{\circ}$ and $t_{\star}$ defined in 3.3, we obtain the following explicit form,
@@
    $\rho_{n}:=\frac{1}{2}\sigma_{n}^{2}+\mathfrak{a}^{n\cdot}\mu+\frac{1}{2}\varsigma^{2}\frac{c_{\Gamma}^{2}}{\lambda_{\Gamma}^{2}}|\mathfrak{a}^{n\cdot}|^{2}|\Sigma|^{2}<r,$ (3.21)
@@
In $(iii)$ of proposition 3.7, $\mathbb{E}\left[\left|\frac{V^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}-\frac{\mathcal{V}^{n}_{t,\mathfrak{d}}}{F^{n}_{t,\mathfrak{d}}}\right|\right]\leq C\varsigma$ means that when the noise term in the productivity process is small, the terms $\frac{\mathcal{V}^{n}_{t}}{F^{n}_{t}}$ and $\frac{V^{n}_{t}}{F^{n}_{t}}$ become closer. The following corollary gives (conditional) laws of the (proxy) of the firm value $\mathcal{V}_{\cdot,\mathfrak{d}}^{n}$.
@@ 3.3 A Credit Risk Model without collateral
In this section, we aim to calculate the probability of default of the previous firm (i.e. operating in an economy that undergoes the climate transition described by both the carbon price and carbon intensities).
As Basel Committee on Banking Supervision, (2017), we introduce four credit risk parameters: the probability of default (PD) measures the default risk associated with each borrower, the exposure at default (EAD) measures the outstanding debt at the time of default, the loss given default (LGD) denotes the expected percentage of EAD that is lost if the debtor defaults, and the effective maturity $T$ represents the duration of the credit. With these four parameters, we can compute the portfolio loss $L$, with some assumptions:
@@
We follow in the present work the second definition of default proposed: a firm default when it is over-indebted, that is in fact the same approach used in the structural credit risk models. We retain this definition in this continuous setting in the same way as Bouveret et al., (2025). Therefore, the over-indebtedness of entity $n$ occurs at time $t$ when the firm value $\mathcal{V}^{n}_{t,\mathfrak{d}}$ falls below a given barrier $\mathcal{D}_{t}^{n}$, related to the net debt, namely on the event $\left\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}$.
However, it should be noted that in a continuous time setting, it can be interesting to work in the Black and Cox, (1976) model. Here, the default event depends on the trajectory of the firm value process $\mathcal{V}$. Therefore, at a given time $t$, the firm defaults if it has been over-indebted at least one time during the period $[0,t]$, that is $\left\{\exists s\in[0,t]\text{ such that }\mathcal{V}^{n}_{s,\mathfrak{d}}<\mathcal{D}^{n}_{s}\right\}$. Thus, the default time is given by $\tau^{n}:=\inf\left\{t\geq 0,\quad\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\right\}$.
@@
therefore $\log{\mathcal{V}^{n}_{\cdot,\mathfrak{d}}}$ is a Gaussian process. However, as Azaıs and Wschebor, (2000) summarizes, the computation of the distribution function of the random variable $\inf_{0\leq s\leq t}\log{\mathcal{V}^{n}_{s,\mathfrak{d}}}$ is by means of a closed formula is known only for a very restricted number of stochastic processes as the Brownian Motion, the Brownian Bridge, the Brownian Motion with a linear drift, and the stationary Gaussian processes with relatively simple with covariance. This is not the case here. Consequently, we extend Merton, (1974) model instead of Black and Cox, (1976) model.
@@
where $\mathcal{K}^{n}(\mathfrak{d},t,T,a,\theta)$ as well as $\mathcal{L}^{n}(t,T)$ are defined in Corollary 3.8 and $\Phi$ the Gaussian cumulative distribution function.
The previous results tell us that the probability of the over-indebtedness depends on parameters specific to:
@@
    the climate transition
@@
    to the company (the contract),
@@
    the economy to which the company belongs to:
@@
Remark 3.11.
Motivated by the uncertainty of the climate transition, a natural extension is to introduce a stochastic carbon price (see, for example, Le Guenedal and Tankov, (2025)). Regardless of whether it is correlated with the productivity process, the expressions of the macroeconomic variables (for example, the output in proposition 3.5) remain unchanged; however, their dynamics do differ. Consequently, primarily due to the form of $v$ defined in (3.13) – whose law is a product and quotient of Gaussian random variables –, we could no longer obtain a closed-form expression for the firm value – because filtration $\mathbb{G}$ would not only be adapted to productivity $\Theta$, but also to the carbon price $\delta$ – and then for the probability of default. Numerical methods are therefore required.
In the next section, we will define and express LGD as a function of some guarantees which are affected by the climate transition. Therefore, by assuming that EAD are deterministic and independent of the carbon price, we will also introduce the expected and unexpected losses.
@@ 4 LGD with stochastic collaterals in continuous time
We are in the same framework as in the previous section, but item 3 of 3.9 is not satisfied anymore, therefore the bank could require from each counterpart $1\leq n\leq N$ a (single) collateral $C^{n}$ to secure its debt. Collateral can take the form of a physical asset such as real estate, business equipment, or inventory, or it can be a financial asset such as invoices, cash, or investments. If a firm is over-indebted at time $t$, we assume that the liquidation ends at $t+a$ with $a\in\mathbb{R}_{+}$ where $a$ is the liquidation delay. Moreover, $k\in[0,1)$ represents the fraction of the collateral used to cover liquidations auctions, as well as other legal and administrative procedures.
@@
The potential loss that would be recorded due to the firm default event is the difference between the debt amount $\mathrm{EAD}_{t}^{n}$ and the amount received after the recovery processes on the time horizon. Consequently, if there is not default ($\mathcal{V}^{n}_{t,\mathfrak{d}}\geq\mathcal{D}^{n}_{t}$) or there is default and the collateral liquidated exceed the exposure ($(1-k)e^{-ra}C^{n}_{t+a}\geq\mathrm{EAD}_{t}^{n}$), the loss is zero, and if there is default and if the exposure exceed the collateral liquidated, the loss is
$\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}-\gamma(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})=(1-\gamma)\left(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a}\right),$ (4.1)
where the constant $r$ is the discount rate. The loss, $L^{N}_{t}$, of the portfolio at time $t$, is in fact, defined as
$L^{N}_{t}:=\sum_{n=1}^{N}(1-\gamma)(\mathrm{EAD}_{t}^{n}-(1-k)e^{-ra}C^{n}_{t+a})_{+}\cdot\mathbf{1}_{\{\mathcal{V}^{n}_{t,\mathfrak{d}}<\mathcal{D}^{n}_{t}\}}.$ (4.2)
@@
$\mathrm{L}^{\mathbb{G},N}_{t}:=\mathbb{E}\left[L^{N}_{t}\middle|\mathcal{G}_{t}\right]=\sum_{n=1}^{N}\mathrm{EAD}_{t}^{n}\cdot\mathrm{LGD}_{t,\mathfrak{d}}^{n}\cdot\mathrm{PD}_{t,\mathfrak{d}}^{n},$ (4.3)
@@
Explicitly, in the above theorem, we assume that our portfolio is perfectly fine grained, so that we can approximate $L_{t}^{N}$ – the portfolio loss – by $L_{t}^{\mathbb{G},N}$ – the conditional expectation of loss given the systemic factor. By construction, the loss given default noted $\mathrm{LGD}$ is the percentage of the total exposure that the bank loses when a over-indebtedness occurs. The literature on LGD modeling is fairly extensive. We can distinguish namely, economic modeling Bastos, (2010); Roncalli, (2020) and stochastic modeling Roncalli, (2020)[Page 193], Chalupka and Kopecsni, (2008). As the definition of $\mathrm{PD}$ does not change compare to what we did in section 3 and as $\mathrm{EAD}$ is given, we will focus on $\mathrm{LGD}$ modeling. We can first remark that $0\leq\mathrm{LGD}_{t,\mathfrak{d}}^{n}\leq 1-\gamma$, then the presence of a collateral necessarily reduces $\mathrm{LGD}$.
Other key quantities for the bank to understand the (dynamics of the) risk in the portfolio are the (expected and unexpected) losses and probability of default conditionally to the (information generated by the) risk factors. Precisely, for a date $t$ and a horizon $T$, a bank computes some risk measures at $t$ of its portfolio maturing at horizon $T$.
@@
    The (conditional) expected loss (EL) – the amount that a bank expects to lose on a credit exposure – reads
@@
    The unexpected loss (UL) – the amount by which potential credit losses might exceed the EL – reads for $\alpha\in(0,1)$,
    $\mathrm{UL}^{N,T}_{t}\!(\alpha):=\mathrm{VaR}^{\alpha,N,T}_{t}-\mathrm{EL}^{N,T}_{t},\quad\text{where}\quad 1-\alpha=\mathbb{P}\left[L_{t+T}^{\mathbb{G},N}\leq\mathrm{VaR}^{\alpha,N,T}_{t}\middle|\mathcal{G}_{t}\right],$ (4.6)
    and the scalar $\mathrm{VaR}^{\alpha,N,T}_{t}$ called value-at-risk.
@@
    The stressed loss (or expected shortfall or ES) – the amount by which potential credit losses might exceed the capital requirement – reads:
@@
Knowing PD, LGD, and EAD is enough to determine the losses. We assume EAD is given (item 2 in 3.9). PD is given by (3.25) in section 3.3. We will now determine LGD when a collateral exists. We focus on two types: a financial asset and a property in housing market. We introduce the two following lemmas which will help later on to set up explicit formulas for LGD. Let $n\in\{1,\ldots,N\}$.
@@
$\rho_{t,T,a}^{n}:=\frac{cv_{t,T,a}^{n}}{\sqrt{\mathcal{L}^{n}(t,T)\overline{\mathcal{L}}^{n}_{t,T+a}}},\quad\text{and}\quad\overline{\omega}_{t,T,a}^{n}:=\frac{\log{\left(u\frac{\mathrm{EAD}_{t+T}^{n}}{(1-k)e^{-ra}}\right)}-\overline{\mathcal{K}}^{n}_{\mathfrak{d},t,T+a}}{\sqrt{\overline{\mathcal{L}}^{n}_{t,T+a}}}.$
@@
However, according to (A.3) in section A.2, $\mathbb{E}[e^{\sigma X}\mathbf{1}_{X\leq x,Y\leq y}]=e^{\frac{1}{2}\sigma^{2}}\Phi_{2}\left(x-\sigma,y-\rho\sigma;\rho\right)$, therefore,